OpenStudy (anonymous):

the cost of your favorite lemonade is $3.84 for 3 gallons. write this as a unit rate.

OpenStudy (texaschic101):

3.84/3 = x/1 (3.84 to 3 gallons = x to 1 gallon) cross multiply (3)(x) = (3.84)(1) can you finish and solve for x